### v3

# MRD: Using Physically Based Differentiable Rendering to Probe Vision Models for 3D Scene Understanding

Abstract.
While deep learning methods have achieved impressive success in many vision benchmarks, it remains difficult to understand and explain the representations and decisions of these models. Though vision models are typically trained on 2D inputs, they are often assumed to develop an implicit representation of the underlying 3D scene (for example, showing tolerance to partial occlusion, or the ability to reason about relative depth). Here, we introduce MRD (metamers rendered differentiably), an approach that uses physically based differentiable rendering to probe vision models’ implicit understanding of generative 3D scene properties, by finding 3D scene parameters that are physically different but produce the same model activation (i.e. are model metamers). Unlike previous pixel-based methods for evaluating model representations, these reconstruction results are always grounded in physical scene descriptions. This means we can, for example, probe a model’s sensitivity to object shape while holding material and lighting constant. As a proof-of-principle, here we assess multiple models in their ability to recover the scene parameters of geometry (shape) and bidirectional reflectance distribution function (material). The results show high similarity in model activation between target and optimized scenes, with varying visual results. Qualitatively, these reconstructions can help investigate the physical scene attributes to which models are sensitive or invariant to. MRD holds promise for advancing our understanding of both computer and human vision, enabling us to answer the question of how physical scene parameters cause changes in model responses.
Figure 1. An overview of metamers rendered differentiably (MRD). Starting from a scene with known parameters $\pi$, we sample positions on the unit sphere, and render the set of ground truth images II. A new scene is initialized from some other state $\pi^{\prime}$ (for example, a sphere shape instead of a dragon). The optimization loop then renders the images at the sampled camera origins and computes the loss between the renders and the ground truth $\mathcal{L}$. We compute the gradient w.r.t. the involved scene parameters and backpropagate, updating the target scene parameters (here, geometry) while holding other parameters (e.g. lighting) constant. This enables targeted probing of a neural network’s understanding of scene properties by separating physical causes and can uncover invariance or even equivalence classes.
1. Introduction
Deep learning has revolutionized pattern recognition from visual input. Image-computable models can now perform many tasks with performance matching or exceeding humans, and their activations can correlate highly with visually-driven responses in primate brains (Doerig et al., 2023). However, it remains difficult to explain how and why these models make the decisions they do. Some work (Alicioglu and Sun, 2022; Zimmermann et al., 2021) probes whether models truly understand scenes, and judge whether these explanations might also provide explanations of visual processing in humans and other animals (Neri, 2022; Funke et al., 2021; Wichmann and Geirhos, 2023).
In this work, we demonstrate how a relatively new technology from computer graphics — physically based differentiable rendering (PBDR) — can be used to evaluate 3D scene understanding in image-computable vision models. PBDR allows the reconstruction of physically plausible 3D scene parameters, such as geometry, camera parameters, and material definitions, via optimization with gradient descent. In contrast to approaches using neural networks for inverse rendering (Mildenhall et al., 2020; Ling et al., 2024; Lopes et al., 2024; Ling et al., 2024; Liang et al., 2024), the PBDR approach is always grounded in the physics of light transport, allowing the physical causes of the image to be separated, decomposed and thereby understood.
Applying PBDR allows one to synthesize new physical scene descriptions which, for example, cause matching model activations to a target scene but are physically different (i.e. are model metamers; (Feather et al., 2023, 2019)). This objective has long been used in the study of human vision, first to support trichromatic color theory (Maxwell, 1857), and then for more general image representations (Freeman and Simoncelli, 2011; Wallis et al., 2016; Balas et al., 2009; Wallis et al., 2019), because it allows the identification of perceptual invariants. Here, we combine PBDR with the metamerism objective to create metamers rendered differentiably (MRD). Rather than trying to infer what the model might understand by interpreting noisy pixel images (e.g. in existing synthesis-based explanation methods), the user can interpret vision neural networks by using model representations to reconstruct specific scene parameters, represented in the physical units of the generating scene. This also opens possibilities to fine-tune existing models on specific scene properties.
To demonstrate the usefulness of PBDR as a tool for model interpretability, we evaluate the implicitly-learned 3D knowledge of vision models trained on 2D images. Vision models are assumed to learn about the underlying 3D structure of scenes, even by training only on 2D images. For example, models trained on 2D scenes can perform well on tasks such as novel view synthesis (Mildenhall et al., 2020), depth estimation (Godard et al., 2017), and 3D object reconstruction (Groueix et al., 2018; Wang et al., 2018). We demonstrate MRD in two example settings: first, we investigate the general problem of “material appearance" by investigating the recovery of surface properties (bidirectional reflectance distribution function, BRDF). Second, we examine the recovery of shape (geometry) in the Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018) metric and in ImageNet trained networks such as ResNet-50 (He et al., 2015) and its shape-bias induced version ResNet-50 SIN trained on a stylized version of ImageNet (Geirhos et al., 2019). We also present results for CLIP (Radford et al., 2021) as a common multi-model embedding backbone.
Our specific contributions are (1) a new method to understand learned visual representations of neural networks by linking their activations to physical environmental properties and efficiently optimizing to find invariants, and (2) results evaluating contemporary vision models using this new method. Because it allows the decomposition of model activations into physical causes, we hope that this method will become an important tool for evaluating learned visual representations.
1.1. Related work
Synthesis-based explanation methods have a long history in vision science and computer vision. The central idea is to understand the representations of image-computable vision models by synthesizing new images via an optimization process that minimizes some objective function. (Other visualization-based explanation methods primarily highlight pixels in the target image that are somehow correlated with a model’s decision (Simonyan et al., 2014; Selvaraju et al., 2017; Ribeiro et al., 2016), but will not be discussed further here). In many previous applications, this involved pixel-based gradient ascent: starting from white-noise images, and iteratively adjusting the pixels. This approach has been used to generate equivalence classes (metamers) for image-computable vision models, leading to significant impact in both computer vision and human vision science (Balas et al., 2009; Portilla and Simoncelli, [n. d.]; Freeman and Simoncelli, 2011; Wallis et al., 2016; Boehm et al., 2025; Broderick et al., 2023; Walton et al., 2021; Balas, 2006; Rosenholtz et al., 2012; Feather et al., 2023, 2019). It is also possible to synthesize images that efficiently discriminate between competing models (Golan et al., 2020; Wang and Simoncelli, 2008; Berardino et al., 2017; Feather et al., 2025; Boehm et al., 2025). Feature visualization and related techniques (Erhan et al., 2009; Mahendran and Vedaldi, 2015; Olah et al., 2017; Nguyen et al., 2019) are also forms of synthesis-based explanation. Here, images are typically generated to maximize the activation of a network layer or node. These methods are argued to provide qualitative insights into network activations and therefore explanations for their decisions, though the usefulness of these methods in providing humans with unique explanatory information has recently been called into question (Zimmermann et al., 2021; Borowski et al., 2021; Geirhos et al., 2024). Another approach has been proposed by Veeravasarapu et al. (Veeravasarapu et al., 2017) by training a GAN using a Bayesian generative model receiving a set of scene parameters and a CNN as a discrimnator to optimize the generative model via Bayesian updates to produce 3D scenes in a probabilistic manner.
To pick examples in which to test our method, we considered both general and more specific image-computable representations. First, general representations (foundation models) are interesting due to their ability to generalize to other visual tasks with minimal further training. Here we use CLIP (Radford et al., 2021) – a multi-modal embedding model aligning textual and visual information – as an example general representation, due to its computational tractability and wide use as a backbone for multi-modal and visual tasks (Poole et al., 2022; Mohammad Khalid et al., 2022; Michel et al., 2022; Sanghi et al., 2022). Second, because some of our evaluations are qualitative and we are interested in perceptual similarity, we also consider perceptual similarity metrics, which are created to try to measure the similarity or discriminability of image pairs for humans (for recent overviews see (Kastryulin et al., 2022; Zhu et al., 2024)). We use LPIPS (Zhang et al., 2018) as a popular example, which uses a VGG backbone and its activations to measure perceptual similarity, with fine-tuning on human similarity judgments. We also test the standalone VGG (Simonyan and Zisserman, 2014), the CNN backbone of LPIPS. Third, previous work identified that common neural networks exhibit a bias towards using texture information to categorize images, whereas humans tended to preferentially categorize by shape (texture vs shape bias; (Geirhos et al., 2019)). This work further showed that training on stylized ImageNet (SIN), in which texture information was decorrelated with shape information via style transfer, increased shape-bias in these networks. We therefore compare the shape reconstruction performance of ImageNet-trained and SIN-trained versions of ResNet50 (ResNet-50-SIN), with the hypothesis that the shape-biased network should perform better at 3D shape reconstruction due to the more disentangled representation of shape and material presumably afforded by SIN-training.
Physically based rendering is an area of computer graphics that seeks to solve physically based light transport often including light-surface or light-volume interactions formulated by the rendering equation ((Kajiya, 1986); see Methods). The rendering equation is a recursive formulation, which results in a high-dimensional integration problem with no analytic solution. Approximating the equation requires Monte Carlo methods alongside physical models accounting for energy conservation, i.e., the strength of the light falls off the farther it travels and the number of times it reflects or refracts from a surface. Differentiable rendering allows the optimization of the physical properties of a scene by differentiating the Monte Carlo integration in the rendering equation. The first approaches differentiating the rasterization pipeline appeared in 2014 (Loper and Black, 2014), allowing the reconstruction of scene parameters such as an object’s shape and BSDF. Third-generation differentiable renderers include Mitsuba3 (Jakob et al., 2022), PyTorch3D (Ravi et al., 2020), nvdiffrast (Laine et al., 2020), and psdr (Zhang et al., 2020). However, unlike PyTorch3D and nvdiffrast, Mitsuba3 and psdr can differentiate through integrators solving physically based rendering (path tracing) while preserving derivatives of the operations, allowing the use of gradient-based optimization methods in PBDR. Actually getting general physically based differentiable rendering to work has a number of technical challenges, caused, for example, by discontinuities with changes in lighting or due to edges, which can interfere with or bias gradient computation. Mitigating these discontinuities and making PBDR computationally tractable are active research areas in the field of differentiable rendering (Nimier-David et al., 2020; Vicini et al., 2021; Nicolet et al., 2021; Zhang et al., 2023, 2024; Jakob, [n. d.]); we build upon these advances here.
2. General Methods
In this section, we introduce the preliminaries for our approach to reconstructing model metamers via differentiable rendering. Our goal is to reconstruct a scene and its parameters (or a subset), given a set of latent representations of ground-truth images. Due to the nature of stochastic gradient descent, the reconstruction will approximate the ground truth, and therefore, will inevitably lead to a different set of scene parameters that are perceived as visually similar by the neural networks used. For the reconstruction, we use six different neural networks, which we categorize into classical convolutional neural networks, perceptual metrics, and, lastly, a pair of modern vision transformers.
In the following, we will briefly recap the methods from the different fields on which our work is based.
Model Metamers
We define model metamers as stimuli that produce the same latent representation, because latent equivalence is the most direct and principled notion of representational indistinguishability in deep networks. Modern models operate entirely on their internal feature representations, and any downstream behavior is fully determined by these latent codes. Inputs that map to the same point (or an equivalence neighborhood) in latent space, therefore, belong to the same representational equivalence class and are functionally indistinguishable to the model. This definition also captures the model’s learned invariances, respects the geometry of its representation, and avoids the limitations of pixel-based or perceptual similarity metrics. By establishing the similarity levels achieved in the baseline reconstruction run, we obtain a criterion for declaring metamers in subsequent experiments: here, we define stimuli whose latent similarity matches this reference level as metameric with respect to the model.
2.1. Physically Based Differentiable Rendering
Physically based rendering models image formation as the evaluation of a light transport operator parameterized by a set of scene parameters $\pi$. Here, we summarise the key existing PBDR methods for readers unfamiliar with this literature. We denote the rendering function by
(1) $f(\pi):\mathbb{R}^{n}\rightarrow\mathbb{R}^{3\times\mathrm{width}\times\mathrm{height}},$
where the output is a color image produced by simulating the propagation of light through the scene.
The Rendering Equation.
Image formation is governed by the rendering equation (RE) (Kajiya, 1986), which recursively expresses the outgoing radiance $L_{o}$ at a surface point $\mathbf{x}$ in direction $\omega_{o}$ as
(2) $L_{o}(\mathbf{x},\omega_{o})$ $=L_{e}(\mathbf{x},\omega_{o})+\int_{S^{2}}f_{s}(\mathbf{x},\omega_{i},\omega_{o})\,L_{i}(\mathbf{x},\omega_{i})\,\cos\theta_{i}\,\mathrm{d}\omega_{i},$
where $L_{e}$ denotes emitted radiance and $L_{i}$ the incident radiance arriving from direction $\omega_{i}$ over the hemisphere $S^{2}$. The BRDF $f_{s}$ describes the directional redistribution of incident irradiance into outgoing radiance. The cosine factor $\cos\theta_{i}=\langle\omega_{i},\mathbf{n}\rangle$ accounts for geometric foreshortening with respect to the surface normal $\mathbf{n}$.
Because the RE is recursive, $L_{i}$ generally depends on further scattering events along the path traced by the incoming ray. Analytic solutions are therefore infeasible for realistic scenes.
Monte Carlo Approximation.
To evaluate the integral in Equation (2), Monte Carlo (MC) integration is employed:
(3) $L_{o}(\mathbf{x},\omega_{o})$ $=L_{e}(\mathbf{x},\omega_{o})+\int_{S^{2}}f_{s}(\mathbf{x},\omega_{i},\omega_{o})\,L_{i}(\mathbf{x},\omega_{i})\,\cos\theta_{i}\,\mathrm{d}\omega_{i}$
(4) $\approx L_{e}(\mathbf{x},\omega_{o})+\frac{1}{N}\sum_{i=1}^{N}\frac{f_{s}(\mathbf{x},X_{i},\omega_{o})\,L_{i}(\mathbf{x},X_{i})\,\cos\theta_{i}}{p(X_{i})},$
where $X_{i}$ are samples drawn from a distribution with density pp. Variance reduction strategies such as multiple importance sampling and stratified sampling improve estimator robustness (Veach, [n. d.]).
Path-Space Formulation.
A more general formulation of light transport is obtained by rewriting the RE in the space of all possible light paths $\Omega$ rather than per-surface interactions. Following Veach (Veach, [n. d.]), the contribution of pixel $j$ can be written as
(5) $I_{j}$ $=\int_{\Omega}f_{j}(\bar{\mathbf{x}})\,\mathrm{d}\mu(\bar{\mathbf{x}}),$
where $\bar{\mathbf{x}}=\mathbf{x}_{0}\mathbf{x}_{1}\mathbf{x}_{2}\cdots\mathbf{x}_{k}$ is a path beginning at a point $\mathbf{x}_{0}$ on a light source, undergoing a sequence of scattering events, and terminating at a sensor location. The path contribution function decomposes as
(6) $f_{j}(\bar{\mathbf{x}})=\underbrace{L_{e}(\mathbf{x}_{0}\rightarrow\mathbf{x}_{1})}_{\text{emission}}\;\underbrace{G(\mathbf{x}_{0}\leftrightarrow\mathbf{x}_{1})}_{\text{geometry + visibility}}\;\underbrace{f_{s}(\mathbf{x}_{0}\rightarrow\mathbf{x}_{1}\rightarrow\mathbf{x}_{2})}_{\text{BSDF}}\;\cdots,$
where the geometry term $G$ includes inverse-square falloff, foreshortening, and visibility via a Dirac-delta that nulls contributions from occluded configurations. The measure $\mu$ encodes the sampling distribution over path space.
Inverse Rendering via Differentiable Light Transport.
Our aim is to estimate scene parameters $\pi$ from observed images $y$ by minimizing a reconstruction loss between $f(\pi)$ and $y$. This requires backpropagating gradients through the light transport simulation, i.e.,
(7) $\pi^{\star}=\arg\min_{\pi}\;\mathcal{L}(f(\pi),y),$
necessitating the derivatives of the rendering equation with respect to $\pi$. Since the outgoing radiance $L_{o}$ depends on scene parameters through geometry, materials, lighting, and visibility, we differentiate Eq. (2) to obtain
(8) $\frac{\mathrm{d}L_{o}}{\mathrm{d}\pi}$ $=\frac{\mathrm{d}L_{e}}{\mathrm{d}\pi}+\frac{\mathrm{d}}{\mathrm{d}\pi}\int_{S^{2}}f_{s}(\mathbf{x},\omega_{i},\omega_{o})\,L_{i}(\mathbf{x},\omega_{i})\,\cos\theta_{i}\,\mathrm{d}\omega_{i}.$
The derivative includes contributions from (1) the BRDF parameters, (2) geometric changes affecting $\mathbf{x}$, normal directions, and thus $\cos\theta_{i}$, (3) derivatives of incident radiance $L_{i}$, and (4) visibility terms, which introduce discontinuities that must be handled via techniques such as soft visibility, edge sampling, or regularization (Jakob et al., 2022; Vicini et al., 2021). Equation (8) forms the foundation of physically based differentiable rendering. It enables gradient-based optimization over complex, physically grounded image formation models and underlies the inverse-rendering procedure used in our experiments.
Handling Visibility Discontinuities.
Differentiating light transport is complicated by visibility discontinuities, which occur at object boundaries and shadow edges. Formally, the derivative of the image with respect to scene parameters $\pi$ decomposes into two components: (1) Interior derivatives, arising from smooth variations of geometry, materials, and illumination along paths whose visibility does not change. (2) Boundary derivatives, arising from changes in ray visibility when geometry crosses occlusion boundaries. These terms dominate near silhouettes and shadow boundaries and are responsible for discontinuous behaviour in the rendering function. While standard path‐wise differentiation techniques handle interior terms, boundary terms require specialized sampling near the visibility manifold. Recent work by Zhang et al. (Zhang et al., 2023) provides a principled treatment of these boundary contributions by explicitly sampling near geometry–visibility boundaries. Their approach yields unbiased estimates of both interior and boundary terms and substantially improves gradient accuracy for inverse rendering tasks.
Following the shape-derivative analysis of Zhang et al. (Zhang et al., 2023), the derivative of a pixel intensity $I_{j}$ with respect to scene parameters $\pi$ decomposes into an interior term and a boundary term:
(9) $\frac{dI_{j}}{d\pi}=\underbrace{\int_{\Omega}\frac{\partial f_{j}(\bar{\mathbf{x}})}{\partial\pi}\,d\mu(\bar{\mathbf{x}})}_{\text{interior term}}+\underbrace{\int_{\partial\Omega(\pi)}f_{j}(\bar{\mathbf{x}})\,(\mathbf{v}\cdot\mathbf{n})\,d\sigma(\bar{\mathbf{x}})}_{\text{boundary term}},$
where $\Omega$ is the set of valid light paths, $\partial\Omega(\pi)$ denotes the geometry visibility boundary (e.g., silhouettes or shadow edges), $\mathbf{n}$ is the boundary normal, and $\mathbf{v}$ is the velocity field induced by the change of scene parameters. The interior term corresponds to smooth variations of geometry, materials, or lighting along paths whose visibility remains unchanged. The boundary term accounts for discontinuities introduced when rays cross visibility boundaries. Zhang et al. show that accurate gradient estimation requires explicitly sampling near $\partial\Omega(\pi)$, yielding unbiased estimates of both terms and significantly improving inverse-rendering stability.
2.2. Neural Networks
In this work, we probe six neural networks to examine how sensitive their internal representations are to shape and material. All neural networks run inference and do not update their weights, i.e., model weights are frozen.
Convolutional Neural Networks.
We begin with two classical convolutional architectures: ResNet50 and ResNet50-SIN (Krizhevsky et al., 2012; Geirhos et al., 2019), where the latter is trained on Stylized ImageNet to increase shape bias. The ResNet architecture relies on residual connections that facilitate gradient flow in deep networks, thereby helping earlier layers avoid the effects of vanishing gradients. These models are widely used for object recognition and as generic feature backbones. For our purposes, we use the ImageNet-1k weights for ResNet50 and the Stylized ImageNet weights for ResNet50-SIN. In both cases, we discard the final classification layer and use the 2048-dimensional penultimate layer as our target latent space. Although earlier layers retain more photometric detail (Gatys et al., 2015; Feather et al., 2023, 2019), our aim is not necessarily to reconstruct the original image, but to reconstruct the conceptual representation encoded in the network’s latent space (we return to this point in the Discussion).
Perceptual Metrics.
A second pair of models is drawn from networks used for perceptual similarity metrics. We evaluate LPIPS (with a VGG backbone) (Zhang et al., 2018) and compare it to the same VGG architecture without the LPIPS perceptual-alignment training. Both metrics aggregate multi-layer VGG features to compute similarity scores, while LPIPS further learns human-aligned weighting from human similarity judgments of image patches.
Vision Transformer.
Finally, we include two recent Vision Transformer (ViT) models – DINOv2 (Caron et al., 2021; Oquab et al., 2024) and CLIP (Radford et al., 2021). These models differ from CNNs by relying primarily on self-attention mechanisms, performing scaled dot-product attention over image patches to build global representations. CLIP learns an aligned representation between modalities (language and visual features).
1:PBDR $f$, parameters $\pi$, loss function $\mathcal{L}$, optimizer $\mathcal{O}$, scene $\mathcal{S}$, transformation $\mathcal{T}$
2:for each epoch $e$ in $\{1,\dots,E\}$ do
3:  for each sensor $v$ in $V$ do
4:   $\pi\leftarrow\mathcal{T}(\pi)$ $\triangleright$ Remesh/clamp scene parameters
5:   $\text{render}_{v}\leftarrow f(\mathcal{S},\pi,v)$ $\triangleright$ Render image given scene, parameters and sensor
6:   $l\leftarrow\mathcal{L}(I_{v},\text{render}_{v})$ $\triangleright$ Compute loss
7:   $l$.backward() $\triangleright$ Compute gradients w.r.t. $\pi$
8:   $\mathcal{O}$.step() $\triangleright$ Backpropagation
9:   $\pi$.update() $\triangleright$ Update scene parameters from $\mathcal{O}$
10:  end for
11:end for
Algorithm 1 The default optimization loop we use to run the experiments. The only factor changing is the transformation function $\mathcal{T}$ we apply to the different experiments, such as remeshing for geometry, or clamping values in the BSDF experiments.
2.3. Scene
Because we are using Mitsuba (Jakob et al., 2022)(see below), we need to define our 3D scene in a Python dictionary or use the XML scene description. In contrast, neural networks do not require an initial scene configuration; we can learn the scene’s assets directly from the images. Therefore, with Mitsuba, we need to initialize a scene that we adjust with gradient updates. In this section, we will introduce the assets and the scene setup.
In this work we use artificial scenes in order to have a well-defined ground truth. We use two separate scenes for shape and material reconstruction.
Shape Reconstruction
The scene consists of multiple environment maps that account for different lighting conditions, and a single floating object, with a diffuse, non-textured BSDF positioned in the center of the scene. We use a floating object for shape reconstructions in order to allow unobstructed views from all camera perspectives.
Material Reconstruction
The scene for the material reconstruction differs from the shape reconstruction. We place an object on a checkerboard and light the scene with a single environment map. The object will start from an initial, textured Principled BSDF (Disney BSDF) (Burley, [n. d.], 2015).
Camera Views
The reconstruction of scene parameters, such as shape and reflectance, benefits from multiple views, which constrain the optimization process. For shape reconstruction, we sample $n$-camera origins using the Fibonacci lattice on a unit sphere, and for material reconstruction, we view the object from front/back and both sides. From these camera positions, we render our $n$-ground truth images and apply the Optix Denoiser (Chaitanya et al., 2017) available in Mitsuba.
Image Format
To ensure that the rendered images produced by Mitsuba 3 are compatible with convolutional neural networks, we convert the output from the renderer’s default high-dynamic-range (HDR) linear radiance representation into a standard display-referred sRGB space. Since the raw Mitsuba outputs contain unbounded scene-referred radiance values that are out-of-distribution for networks such as ResNet, we first transform the renderer’s internal RGB basis to linear sRGB, and then apply Reinhard tonemapping to compress the dynamic range into a bounded interval. Reinhard tonemapping compresses unbounded HDR radiance into a bounded, perceptually plausible range by applying a simple luminance normalization that preserves detail while preventing saturation. Finally, we apply the sRGB transfer function (gamma encoding), yielding images with normalized, perceptually uniform intensities suitable for training. This conversion ensures consistent color representation and prevents numerical instabilities arising from unbounded HDR values during learning.
2.4. Implementation Details
We implement all of our experiments using Python 3.12, PyTorch (Paszke et al., 2019), and Mitsuba 3 (Jakob et al., 2022), which is a suitable PBDR for our experiments, and mitigates many of the discontinuity issues mentioned above ((Nimier-David et al., 2020; Vicini et al., 2021; Nicolet et al., 2021; Zhang et al., 2023, 2024)). For all of our experiments, we also define an early stopping criterion: the experiment stops if the loss does not improve for 50 epochs. We run all of our shape reconstruction experiments on a workstation with a single RTX 4090 with 24 GB of VRAM using Ubuntu 22.04.5 LTS. For our material reconstruction experiments, more VRAM is required due to the path tracing integrator; for these experiments, we resort to an RTX 6000 with 48 GB of VRAM running on an identical system environment. We use pretrained models and implementations of these if available (Geirhos et al., 2019; Radford et al., 2021). Our code will be available here: https://github.com/ag-perception-wallis-lab/MRD.
2.5. Baseline experiments
To calibrate our optimization processes and ensure that ground truth reconstructions can be obtained external to any learned representation, we implement baseline experiments for both shape and BSDF reconstruction. We require the baselines to reconstruct the scene parameters we are optimizing. Multiple scene parameters affect the reconstruction fidelity, such as the samples per pixel, the number of views, and the integrator used to solve the rendering equation.
The baselines use pixel-based losses. For shape, that is mean-absolute error, which works overall because it is more robust to outliers. For translucent material reconstruction, we have to use the path tracing integrator within Mitsuba (Jakob et al., 2022). This integrator has a substantially larger memory footprint than the other integrators present in Mitsuba. However, recent integrators do not allow differentiation of parameters that affect transmittance, therefore translucency. Deng et al. introduce the Dual-Buffer loss (Deng et al., 2022), which exhibits good convergence properties while requiring fewer samples per pixel, making path tracing’s usage possible. The Dual-Buffer loss computes the loss between two rendered images $\hat{y},\hat{y}\prime$ and the ground truth view $y$,
(10) $\mathcal{L}_{dual}(\hat{y},\hat{y}\prime,y)$ $=\frac{1}{N}\sum_{i=0}^{N}\left((\hat{y}-y)\cdot(\hat{y}\prime-y)\right)^{2}.$
Whenever we use the path tracing integrator for our materials, we also use the Dual-Buffer loss. For shape baselines, we use mean-absolute error.
The general optimization loop is displayed in Algorithm 1, and the overview of the hyperparameters for each scene is shown in Tables [1,2].
Table 1. Hyperparameters for material reconstruction.
Material # Views LR Epochs BSDF Parameters
Translucent 4 $2\times 10^{-1}$ 500 roughness, $\eta$, albedo texture, spec_trans
Diffuse 4 $3\times 10^{-2}$ 500 all (albedo texture)
Brushed Metal 4 $3\times 10^{-2}$ 500 all (albedo & anisotropic texture)
Aurora 4 $3\times 10^{-2}$ 500 all (albedo texture)
Table 2. Hyperparameters for shape reconstruction.
Shape # Views $\boldsymbol{\lambda_{\text{reg}}}$ LR Epochs Remesh Epochs
Dragon 25 15 $1\times 10^{-1}$ 500 [5, 25, 50, 100, 150, 250, 350, 450]
Lion Statue 25 15 $1\times 10^{-1}$ 500 [5, 25, 50, 100, 150, 250, 350, 450]
Dog 25 15 $1\times 10^{-1}$ 500 [5, 50, 100, 150, 250, 350, 450]
Suzanne 8 25 $1\times 10^{-1}$ 500 [100, 200, 300, 400]
2.6. Analysis
In this subsection, we will cover the techniques we use to quantify our results. We define two metrics, unit hypersphere similarity and representational similarity analysis, between those latent representations.
Using similarity on the unit hypersphere allows us to factor out variations in vector magnitude and focus exclusively on directional information in the latent representations. This is particularly beneficial in our setting, since many latent spaces—especially those produced by deep networks—encode semantic content primarily through the orientation of latent vectors rather than their norm, which can be affected by scale, activation statistics, or training dynamics. By normalizing all representations, we obtain a well-behaved similarity measure (cosine similarity) that directly reflects angular agreement between predicted and ground-truth codes, thereby providing a scale-invariant and geometry-preserving comparison. Complementarily, RSA serves as a higher-level measure that evaluates not only pointwise similarity between individual latent codes but also the relational structure induced over an entire set of views. This global perspective is crucial when reconstructing a latent space: even if individual codes align in isolation, the reconstruction is only meaningful if the pairwise similarity structure, the geometry of the manifold, is preserved. RSA therefore provides a principled way to assess whether the reconstructed latent space captures the same representational organization as the ground-truth latent space, thereby complementing hyperspherical similarity.
Similarity on the Unit Hypersphere.
Let $z\in\mathbb{R}^{d}$ denote a latent representation. We map $z$ onto the unit hypersphere by $\ell_{2}$-normalization:
(11) $\phi:\mathbb{R}^{d}\to\mathbb{S}^{d-1},\qquad\phi(z)=\frac{z}{\lVert z\rVert_{2}+\varepsilon},$
where $\varepsilon>0$ is a small constant added for numerical stability, and $\mathbb{S}^{d-1}=\{x\in\mathbb{R}^{d}:\lVert x\rVert_{2}=1\}$ denotes the unit hypersphere in $\mathbb{R}^{d}$.
Given two latent vectors $z_{\text{render}},z_{\text{target}}\in\mathbb{R}^{d}$, we first normalize them:
(12) $\hat{r}=\phi\bigl(z_{\text{render}}\bigr)=\frac{z_{\text{render}}}{\lVert z_{\text{render}}\rVert_{2}+\varepsilon},\qquad\hat{t}=\phi\bigl(z_{\text{target}}\bigr)=\frac{z_{\text{target}}}{\lVert z_{\text{target}}\rVert_{2}+\varepsilon}.$
The similarity score used in our method is then the inner product between the two normalized vectors:
(13) $\mathrm{sim}\;=\;\hat{r}^{\top}\hat{t}$.
Since $\lVert\hat{r}\rVert_{2}=\lVert\hat{t}\rVert_{2}=1$ by construction, this inner product corresponds to the cosine of the angle $\theta$ between them:
[…]
Representational Similarity Analysis (RSA)
Let $\{x_{1},\dots,x_{N}\}$ denote a set of $N$ views (images). We consider two latent representations of these views, for example, produced by two different scenes:
$z_{i}^{(1)}=g_{1}(x_{i})\in\mathbb{R}^{d_{1}},\qquad z_{i}^{(2)}=g_{2}(x_{i})\in\mathbb{R}^{d_{2}},\quad i=1,\dots,N.$
We first map each latent vector onto the unit hypersphere (as in our similarity measure above) using $\ell_{2}$-normalization
$\hat{z}_{i}^{(m)}=\frac{z_{i}^{(m)}}{\lVert z_{i}^{(m)}\rVert_{2}+\varepsilon},\qquad m\in\{1,2\}.$
For each representational space $m$, we then construct a representational similarity matrix (RSM) $S^{(m)}\in\mathbb{R}^{N\times N}$ with entries
(15) $S^{(m)}_{ij}\;=\;k\!\bigl(\hat{z}_{i}^{(m)},\hat{z}_{j}^{(m)}\bigr),\qquad k(a,b)=a^{\top}b,\qquad 1\leq i,j\leq N,$
i.e. the cosine similarity between the two normalized latent vectors. (Equivalently one may define a dissimilarity matrix $D^{(m)}_{ij}=1-S^{(m)}_{ij}$.)
RSA quantifies how similar the geometry of these two representational spaces is, by correlating their RSMs. Let $\operatorname{vec}_{\triangle}(\cdot)$ denote vectorization of the upper triangular (off-diagonal) entries of a matrix. The RSA score is then defined as
(16) $\rho_{\mathrm{RSA}}\;=\;\mathrm{corr}\!\left(\operatorname{vec}_{\triangle}\bigl(S^{(1)}\bigr),\operatorname{vec}_{\triangle}\bigl(S^{(2)}\bigr)\right),$
where $\mathrm{corr}$ is a Kendall rank correlation coefficient. A high value of $\rho_{\mathrm{RSA}}$ indicates that pairs of views that are similar (or dissimilar) in one latent space tend to be similarly related in the other latent space, i.e. the two spaces encode a similar representational structure over the set of views.
Kendall’s $\tau$, as implemented in scipy.stats.kendalltau, measures the strength of a monotonic relationship between two variables by comparing the number of concordant and discordant pairs of observations. SciPy computes the $\tau$-b variant, which corrects for ties in both variables. The test statistic is then evaluated for significance using either an exact permutation distribution (for very small sample sizes without ties) or, more commonly, a large-sample normal approximation. In the latter case, SciPy computes the variance of $\tau$ under the null hypothesis of independence and converts the test statistic to a z-score, from which the two-sided p-value is obtained. We consider our experiments significant if the hypothesis test’s $p<0.05$.
Figure 2. Top: Hyperspherical similarity between reconstructed and ground-truth latents across optimization, shown for all feature representations. These curves define the expected similarity levels obtainable when the BSDF is known, providing the reference with which later, unknown-material reconstructions are compared. Shaded regions show one standard deviation across views. Middle: Reconstructed images at selected iterations (5, 15, 25, 175); each image shows one of the four object views used in these experiments. As similarity increases, the material appearance—including specular structure and anisotropic reflections—becomes increasingly consistent with the ground truth. Bottom: Corresponding error visualizations (warm yellow/red = higher error, cool blue lower error) highlight how discrepancies contract as the optimization progresses, eventually capturing the reflectance structure characteristic of the brushed-metal BSDF. These results serve as the calibration signal for determining whether subsequent BSDF-based reconstructions achieve metameric similarity.
3. Material Reconstruction
Every object in our scene has one or multiple materials attached to it that define how light reflects and/or refracts when light interacts with the surface or media. It is one of the main drivers of how light is transported through space.
In our BSDF optimization experiments, we make parameters of a BSDF model available as targets for optimization, fixing other scene parameters (geometry, lighting, etc). Our optimization then aims to reconstruct the BSDF (partially with textured BSDFs). The difficulty in this task lies in the absence of a one-to-one parameter mapping, as the Principled BSDF has no direct parameter relationship to the other materials and is therefore metameric by definition. This forces the optimization process to choose from a set of parameters to replicate the models’ representation of the target material. Because this run uses the ground truth pixel loss as optimization target, the resulting similarity trajectories reveal the level of agreement that can be achieved when the reconstruction problem is well-posed and no model mismatch is present. This baseline is essential for interpreting later experiments: we define a successful reconstruction, i.e., the reconstruction of the scene as a model metamer, if it achieves similarity values comparable to those obtained in this baseline setting.
3.1. Methods
Principled BSDF
The Principled BSDF (Burley, [n. d.], 2015) was introduced to provide a production-friendly, art-directable surface model that remains physically plausible while exposing an intuitive parameter set. Rather than representing a single physical model, the Principled BSDF blends a small set of carefully chosen analytic lobes – diffuse, retro-reflective, specular microfacet, clearcoat, sheen, and transmission – into a unified framework with consistent energy conservation. Each lobe is designed to behave reasonably across the full parameter domain, enabling smooth transitions between dielectric and metallic appearances via a single metalness control while preserving reciprocity and approximate Fresnel behavior (which also benefits the optimization due to predictable changes).
Integrator
Reconstructing a Principled BSDF in Mitsuba requires the use of a path tracing integrator because the appearance of such materials is determined by a combination of diffuse, specular, microfacet, subsurface, and transmission components that interact nonlinearly with the full distribution of incoming light. Translucent effects cannot be captured using runtime cost-efficient integrators such as Path Replay Backpropagation (Vicini et al., 2021). Path tracing provides an unbiased estimate of the rendering equation, allowing gradients to flow through all physically relevant light transport paths. This is essential for recovering BSDF parameters whose influence may only emerge through indirect bounces, grazing-angle reflections, or jointly through coupled terms such as roughness, index of refraction, and anisotropy. Consequently, a physically based global illumination integrator like path tracing is necessary to ensure that the optimization receives accurate, informative feedback and can faithfully reconstruct the full behavior of a Principled BSDF. Further, the path tracing integrator requires more samples per pixel to reconstruct the textures used to parameterize the BSDF.
Parameter Clipping
To ensure physically plausible behavior during optimization, the parameters of the Principled BSDF in Mitsuba must be clamped to their valid ranges. Many of these parameters, such as roughness, specular transmission, metalness, and anisotropy, are defined only over restricted intervals $[0,1]$, and allowing the optimizer to explore values outside these domains can lead to numerical instabilities, non-physical reflectance, or invalid configurations within the renderer. Clamping prevents such degeneracies by projecting intermediate updates back into the feasible set, ensuring that the material model remains consistent with its underlying physical assumptions throughout the optimization process. This constraint is particularly important when gradients are noisy, as is common in path tracing, where unconstrained parameter updates may otherwise accumulate error and destabilize convergence. By enforcing these bounds, we maintain stable optimization dynamics and guarantee that the reconstructed material remains compliant with the Principled BSDF specification.
3.2. Results
Figure 2 shows the baseline optimization and model similarity scores for one of our scenes (brushed-metal). Because the optimization in this setting targets the ground truth pixel-based loss, the resulting similarity curves reveal how closely ideal conditions produce high similarity scores in each latent representation. All networks exhibit a consistent increase toward their characteristic asymptotic similarity: early iterations rapidly capture the dominant reflectance cues, while later iterations refine finer-scale material structure, such as anisotropy and shading variations. The qualitative reconstructions confirm this progression, demonstrating that as latent similarity approaches the ground-truth baseline, the rendered images converge toward the correct brushed-metal appearance. This baseline, therefore, serves as the material-specific metamer criterion; similar baselines were created for other materials. The baseline run is therefore not simply a diagnostic check but a calibration experiment. It establishes the similarity levels that correspond to true metameric alignment in each representation. The subsequent experimental results can therefore be interpreted relative to these baselines: if a method achieves ground-truth-level similarity for a given network, we conclude that it has produced model metamers with respect to that network’s representational geometry.
The reconstruction curves in Figure 3 show that both VGG and ResNet rapidly converge toward the latent representation characteristic of the brushed-metal material, though the networks reach the baseline at different rates and with different asymptotic behaviors. VGG approaches the baseline more slowly and exhibits a modest but persistent gap, consistent with its lower representational precision for this material class. In contrast, ResNet closely tracks the baseline curve and reaches the baseline similarity threshold early in optimization, indicating a high degree of latent alignment with the target brushed-metal structure. The accompanying renderings corroborate these tendencies: early iterations capture coarse geometry and broad reflectance structure, while later iterations refine specular highlights, anisotropic streaking, and subtle shading cues associated with brushed metal.
Figure 3. Reconstruction similarity compared to the ground-truth baseline for brushed-metal objects using VGG (top) and ResNet (bottom). Left: Hyperspherical similarity between the reconstructed latent code and the target latent representation over optimization iterations (blue), shown alongside the baseline similarity (orange) obtained during the ground-truth run for this material. The horizontal red marker on the ResNet curve indicates the iteration at which reconstruction first intersects the baseline similarity threshold. Right: Corresponding renderings at selected iterations illustrate the progressive refinement of material appearance, compared with the ground-truth baseline images.
Table 3. Material reconstruction results across all material models and NNs. In addition to the experiment label, we show the epoch of peak similarity, the peak similarity itself, the corresponding baseline value, the RSA (when applicable), the $p$-value for the RSA correlation (see text) and the signed similarity difference.
Experiment (BSDF + Model) Epoch Top Sim Top Base RSA Significant $(p<0.05)$ Similarity Difference
METAMERS Aurora CLIP 64 0.883 0.848 1.000 Yes 0.035
Aurora DINO 127 0.955 0.914 0.733 No 0.041
Aurora LPIPS 24 0.969 0.958 - - 0.011
Aurora ResNet-SIN 45 0.999 0.999 0.2 No 0.000
Aurora VGG 25 0.932 0.910 - - 0.022
Brushed Metal CLIP 122 0.756 0.731 -0.067 No 0.025
Brushed Metal DINO 114 0.811 0.784 0.600 No 0.027
Brushed Metal LPIPS 197 0.920 0.891 - - 0.029
Brushed Metal ResNet-SIN 110 0.998 0.998 0.333 No 0.000
Brushed Metal VGG 199 0.842 0.771 - - 0.071
Diffuse DINO 199 0.795 0.790 0.467 No 0.006
Diffuse LPIPS 199 0.884 0.851 - - 0.033
Diffuse ResNet-SIN 130 0.998 0.998 0.467 No 0.000
Diffuse VGG 199 0.793 0.690 - - 0.103
[…]
Across our different scenes, we were able to find metamers for most of the networks tested (Table 3). The translucent reconstructions, ResNet and CLIP, had examples that did not qualify as metamers under our definition. Subjective visual results appeared reasonable (Figure 3 shows example reconstructions of the brushed metal materials for VGG and ResNet).
Summary of Material Reconstruction Outcomes
Table 3 summarizes the per-BSDF / per-network outcomes. The metameric results attain a top similarity that meets or exceeds the baseline (similarity difference $\gtrsim 0$), indicating that the reconstructed BSDF is representationally indistinguishable from the ground truth under the evaluated representation. In the case of non-metamers, the top similarity is noticeably lower than the baseline (negative similarity difference), indicating that the reconstruction did not satisfy the material metamer criterion for that network. Overall, we find 14 metamers and 10 non-metamers.
Representation-dependent Behaviour
The success rate depends on the choice of feature representation. Perceptual metrics and scale-invariant CNN features (e.g. LPIPS, ResNet-SIN) frequently reach baseline-level similarity (note multiple LPIPS and ResNet-SIN rows in the metamers with vanishing differences), whereas some self-supervised and multimodal networks (e.g. certain CLIP and DINO entries) show larger variability. ResNet-SIN in particular repeatedly attains near-identical peak and baseline similarities (differences $\approx 0$), suggesting that this representation aligns especially well with the BSDF reconstruction objective. VGG and LPIPS also perform reliably for many BSDFs, but VGG exhibits larger gaps in a few cases (e.g. diffuse and brushed-metal entries), consistent with a coarser or less material-sensitive encoding in those instances.
BSDF-specific Effects and Difficulty Regimes
Different material classes show systematically different recoverability. Metallic and specular BSDFs (brushed-metal, aurora-like materials) tend to produce strong global image cues (high-contrast highlights, anisotropic streaks) that generate coherent gradients and are thus easier to recover to baseline similarity. Diffuse and translucent materials can be more challenging: they sometimes produce broader similarity gaps or lower RSA values, likely because their appearance depends on subtle indirect-lighting interactions or transmission effects that are harder to disambiguate from other scene factors. Where translucency interacts with geometry and indirect illumination, the optimization needs a larger number of samples and iterations to reach the same representational fidelity.
RSA, Pointwise vs. Relational Alignment
We compute RSA for NNs producing a latent representation used for optimization, i.e., ResNets, DINOv2 and CLIP. RSA values in Table 3 provide a complementary view on success. In several cases where peak similarity is high but RSA is modest or non-significant, reconstructions match the ground-truth latent vector for the target view (pointwise alignment) but do not fully reproduce the global pairwise geometry of the representation across views. Conversely, positive and significant RSA together with high peak similarity indicate both pointwise and manifold-level agreement. The generally mixed RSA values suggest that for some BSDFs, the optimization recovers correct per-view latents without completely restoring the representational geometry; this may be due to residual view-dependent effects or slight mismatches in how different views sample highlight and grazing-angle behaviour.
Noise, Clipping and other Practical Considerations
Material reconstruction relies on unbiased path tracing and, therefore, is subject to Monte Carlo noise in both forward rendering and gradient estimates. We found that (1) parameter clipping to the physically valid domain is essential to avoid unstable and physically implausible updates, and (2) increasing sample count can reduce noisy gradients and improve final alignment. (3) For complex scenes, truncating the maximum number of paths per ray can also affect final results, especially for translucent material, where light can refract and scatter multiple times within the same object.
Implications
Together, these results reinforce that metamericity is both representation- and target-dependent. Materials generally admit tighter, lower-dimensional manifolds in representation space and therefore are more readily recovered to baseline-level similarity than high-dimensional shape degrees of freedom (see below). Nonetheless, not all representations treat materials the same: some networks encode subtle BSDF-dependent cues more faithfully than others.
4. Shape Reconstruction
Geometry in computer graphics is often represented as a triangulated mesh using vertices and edges to form faces, and therefore, the surface of an object. A defined set of vertices defines a single face; however, the vertices can still move in the three-dimensional space and form wider or narrower faces/triangles. In the shape reconstruction experiments, we allow moving the vertices in space to model the target shape visible in the reference images. In our experiments, we start from a proxy mesh – an icosphere with 16000 vertices – optimizing it towards the target mesh.
4.1. Methods
Large Steps
Our implementation for the shape reconstruction follows the Large Steps method of Nicolet et al. (Nicolet et al., 2021). This method achieves fast and efficient convergence by preconditioning the bias of the gradient steps towards smooth solutions without enforcing smoothness in the final result. $\lambda$ is a new hyperparameter introduced in (Nicolet et al., 2021) that controls the strength of regularization introduced by the Laplacian matrix. Nicolet et al also introduce changes in the Adam (Kingma and Ba, 2017) optimizer to precondition the gradient update and allow for higher learning rates. In our implementation, we use Mitsuba 3 instead of nvdiffrast and PyTorch, necessitating adjustment of $\lambda$ and the step size, thus deviating from the suggestions in (Nicolet et al., 2021).
Integrator
We use the direct projective path replay backpropagation method introduced by Zhang et al. (Zhang et al., 2023) for our shape reconstruction experiments. This integrator is specifically designed to handle the discontinuities that arise when object boundaries move during optimization. It mitigates these boundary discontinuities by using projective sampling, where samples generated during the forward pass are reprojected onto nearby geometry boundaries to estimate the boundary derivative stably and with reduced variance. The key contribution of Zhang et al. is to rewrite the boundary term of the rendering equation into a local integral over projected samples, removing the need to integrate over the full non-local geometry. This projective formulation offers several advantages: (1) it eliminates the complex non-local domain that normally appears in boundary derivatives, (2) it removes the geometric term $G$, which is responsible for much of the variance in classical formulations, and (3) it leads to a significantly simplified expression for common geometries such as smooth closed surfaces and polygonal meshes. A helpful way to build intuition is to consider a simplified 2D example: imagine a curtain in front of a window. Moving the curtain changes the visibility discontinuously. Instead of tracking how the curtain itself moves (which produces the discontinuity), the integrator conceptually “moves" the image of the window onto the curtain, i.e., it adjusts the integration domain rather than the geometry. This projection makes the derivative well-behaved and substantially reduces estimator variance.
Tesselation
The target mesh can use more vertices than our initial icosphere. If that is the case, we allow our current mesh reconstruction to tessellate its mesh using each side’s midpoint (Botsch and Kobbelt, 2004). We apply this remeshing operation on specific epochs, which we pick after testing our baseline. Tesselating the mesh will always result in an increase in the loss, because we converged to the vertices’ position, yet adding more vertices allows the optimization to deal with details on the object’s surface. Each tesselation will be followed with a learning rate decay multiplying the previous learning rate by a factor of $8e^{-3}$. We apply this decay only after remeshing because the details in the mesh no longer require large changes.
Figure 4. Top: Hyperspherical similarity/LPIPS between reconstructed latents and ground-truth latent codes over the course of training. This run serves as a diagnostic reference, illustrating how rapidly and how well different network feature spaces can be aligned under our chosen optimization settings. Shaded regions show one standard deviation across views. Middle: Reconstructed images at selected iterations (5, 25, 75, 446) during the baseline run (grid of images show the 25 views). As training progresses, reconstructions transition from coarse structural estimates to high-fidelity matches of the target views, demonstrating the optimizer’s capability in the idealized setting. Bottom: Corresponding error/activation visualizations for the same iterations. Early iterations show widespread, high-magnitude discrepancies, which gradually contract into localized residuals as the reconstructed latent space aligns with the ground truth. This behavior validates the optimization dynamics used for all subsequent experiments.
4.2. Results
The results in Fig. 4 summarize the behavior of our method during the baseline (mean-absolute error) reconstruction run, whose primary role is to establish a representative similarity baseline for all subsequent experiments (as in the Material reconstructions above).
Across networks, we observe a consistent pattern: a rapid initial rise in similarity followed by gradual convergence. Importantly, even though the absolute similarity varies across architectures, each network reaches a characteristic asymptotic similarity value that reflects how tightly the reconstructed geometries can align with their true directions under ideal conditions. These asymptotic levels thus serve as the expected similarity signatures for each feature space. When later experiments achieve similarities in this range, we can meaningfully claim that they produce model metamers according to that representation.
The qualitative reconstructions in the baseline run reinforce this interpretation. As training progresses, the renderings evolve toward highly faithful reproductions of the target images. The tightening of error maps mirrors this improvement, indicating that the optimizer successfully recovers both the coarse and fine structure encoded in the ground-truth representation. This process provides a visual intuition for what it means to “match” a latent code: when latent similarity approaches the ground-truth reference level, the reconstructions are representationally indistinguishable from the originals (while remaining physically different).
Figure 5 demonstrates the convergence behavior and resulting metameric geometries for two of our experiments (Dragon Hallstatt using LPIPS as the metric / latent space, and Dragon Garden with ResNet as the latent). The reconstruction similarity consistently improves, converging towards the baseline performance (orange curve). We specifically analyze the iteration where the reconstruction fidelity peaks or intersects with the baseline threshold. In the Dragon Hallstatt LPIPS experiment, the reconstruction similarity exceeds the baseline similarity (making this a metamer according to our definition). The visual output at this intersection point is subjectively indistinguishable from the baseline. However, visual results for the Dragon Garden ResNet experiment are strikingly different. First, the reconstruction similarity does not exceed the baseline similarity, making this non-metameric by our strict definition. However, the two similarities are quite close, and in addition the RSA correlation of 0.67 is significant (see Table 4). Visually, however, the ResNet results look quite unlike a dragon to a human viewer. That the similarities and RSA scores for ResNet are high despite these qualitative differences is consistent with the notion that ResNet has a very wide equivalence class for 3D shapes: spiky blobs look like dragons to ResNet. We observed similar results for ResNet applied to other geometries.
Figure 5. Evaluation of reconstruction fidelity using LPIPS similarity and ResNet hypersphere similarity. We compare the convergence of our scene parameter reconstruction (blue) against the baseline similarity (orange) for the Dragon Hallstatt (top) and Dragon Garden (bottom) scenes. The plots track the perceptual similarity over iterations, with the visual results displayed at the point of peak similarity (marked by the intersection, no intersection $\rightarrow$ model metamer). The side-by-side comparisons demonstrate that when the reconstruction similarity aligns with the baseline, the resulting synthesized views are perceptually equivalent to the model references.
Table 4. Shape reconstruction results across all shapes, environment maps and NNs. The table shows the epoch of the top recorded similarity (averaged across views) and compares it to the top similarity measured during base reconstruction. If the NN produces a latent space we further measure the RSA and its significance values. Furthermore, the similarity difference is provided to see the marginal errors. Lastly, we compute the difference between seeds and baseline/null ECDFs (see Figure 6, approximate a Gaussian kernel density and normalizing the $x$-range to [1,1][-1,1]. The vertical bars show [1,0,1][-1,0,1] from left to right. This table is best viewed digitally.
Experiment (Shape + Envmap + Model) Epoch Top Sim Top Base RSA Significant ($p<0.05$) Similarity Difference Seed Null Diff Seed Base Diff
METAMERS Dog Garden LPIPS 159 0.991 0.983 - - 0.008 [Uncaptioned image] [Uncaptioned image]
Dog Garden VGG 164 0.983 0.963 - - 0.020 [Uncaptioned image] [Uncaptioned image]
Dog Hallstatt LPIPS 158 0.991 0.991 - - 0.000 [Uncaptioned image] [Uncaptioned image]
Dog Hallstatt VGG 258 0.982 0.979 - - 0.003 [Uncaptioned image] [Uncaptioned image]
Dog Skybox VGG 499 0.980 0.835 - - 0.144 [Uncaptioned image] [Uncaptioned image]
Dragon Skybox LPIPS 248 0.798 0.782 - - 0.016 [Uncaptioned image] [Uncaptioned image]
Dragon Skybox ResNet 238 0.953 0.945 0.536 Yes 0.008 [Uncaptioned image] [Uncaptioned image]
Dragon Skybox VGG 486 0.578 0.519 - - 0.058 [Uncaptioned image] [Uncaptioned image]
Lion Statue Hallstatt LPIPS 348 0.957 0.956 - - 0.001 [Uncaptioned image] [Uncaptioned image]
Suzanne Garden LPIPS 115 0.986 0.972 - - 0.015 [Uncaptioned image] [Uncaptioned image]
Suzanne Garden VGG 118 0.970 0.937 - - 0.034 [Uncaptioned image] [Uncaptioned image]
Suzanne Hallstatt LPIPS 159 0.985 0.979 - - 0.006 [Uncaptioned image] [Uncaptioned image]
Suzanne Hallstatt VGG 101 0.968 0.953 - - 0.015 [Uncaptioned image] [Uncaptioned image]
Suzanne Skybox LPIPS 236 0.953 0.933 - - 0.020 [Uncaptioned image] [Uncaptioned image]
Suzanne Skybox VGG 158 0.872 0.849 - - 0.023 [Uncaptioned image] [Uncaptioned image]
[…]
Summary of Shape Reconstruction Outcomes
Table 4 reports, for each experiment, the epoch at which the reconstruction attained its peak (averaged) similarity, the corresponding peak similarity, the baseline peak similarity, an RSA measurement where applicable, and the signed difference between peak and baseline similarities. In the metameric set of cases, the reconstruction attains a top similarity that meets or slightly exceeds the baseline similarity (similarity difference $\approx 0$ or positive), indicating that the reconstructed scene is representationally indistinguishable from the ground truth under the evaluated representation. In the other and larger set of cases, the top similarity falls noticeably short of the baseline (negative similarity difference), indicating failure to reach the baseline metamer criterion for that particular representation. By our definition, we find 15 model metamers and 56 non-metamers.
Network-dependent Behaviour
The results exhibit a clear dependence on the choice of feature representation. The LPIPS perceptual metric and the VGG convolutional backbone frequently reach baseline-level similarity (several LPIPS and VGG rows appear in the metameric set with very small differences, e.g. Dog Hallstatt LPIPS with only a small difference), whereas several self-supervised or multimodal representations (for example many CLIP and DINO runs) tend to produce larger negative gaps and thus are more often classified as non-metamers. ResNet-SIN commonly attains very high peak similarities (often within $10^{-3}$ of the baseline), suggesting that the scale-invariant features in this architecture align particularly well with the reconstruction objective. VGG shows mixed behaviour: some VGG runs reach metamer-level similarity while others (especially under more challenging environment maps) stop short.
Environment and Shape Effects
Environment lighting interacts with shape to affect recoverability. For example, the same shape in different environment maps can determine metamerism (compare Dog Skybox and Dog Garden rows). In general, scenes with complex environment-driven reflections or strong view-dependent shading (e.g. skybox illumination) are more likely to yield larger similarity gaps, consistent with the intuition that view-dependent effects increase the effective complexity of the inverse problem.
RSA; Pointwise and Relational Alignment
Where RSA is reported, it provides a complementary perspective on representational alignment. Several non-metamer cases nonetheless exhibit moderate-to-high RSA values (for example, entries with large negative similarity differences but RSA values >0.5>0.5), indicating that while the reconstruction may not match the ground truth latent vector exactly (pointwise mismatch), it nevertheless preserves the pairwise geometry of the representation (relational agreement). Conversely, low RSA together with a negative similarity difference signals both pointwise and geometric mismatch. Thus, RSA helps disambiguate whether failures are failures of pointwise code recovery or failures of recovering the representational geometry itself.
Figure 6. The mean empirical cumulative distributions of across all experiments (three environment maps, four geometries and six models). We show three distributions. (a) The null distribution (orange) obtained from random, unrelated objects; (b) the seeds distribution (teal) showing variability across independent reconstruction runs with different initializations; (c) and the baseline distribution (purple), representing the similarity achieved during the ground-truth optimization run used for hyperparameter selection. Higher similarity values denote closer alignment between latent representations. The shaded region is one standard deviation.
Control experiments
As a control experiment to verify our criterion for metamerism in the shape experiments, we compute the empirical cumulative distribution function (ECDF) for three shape conditions: (1) a null distribution, being the hyperspherical similarity between the ground truth target mesh renderings and 1000 random object meshes, (2) multiple reconstructions with different seeds, and (3) similarity measured during ground truth run. We show the average ECDF across all shape reconstructions in Figure 6, showing a clear separation of the null and the other distributions. Furthermore, we plot the differences between the null and seed distribution, as well as the difference between null and baseline distribution in Table 4. The null distribution occupies the lowest range of similarities, confirming that unrelated objects produce widely scattered and generally low latent correlations. In contrast, the seeds distribution clusters tightly near high similarity values, indicating that independent reconstruction runs converge to nearly identical similarity values despite differing initial seeds. Finally, the baseline distribution lies at the extreme upper end of the similarity spectrum, reflecting the latent alignment achieved during the ground-truth run. This hierarchical ordering, $null<seeds\approx baseline$, supports our metamer criterion: reconstructions qualify as model metamers when their latent similarity matches the characteristic similarity levels measured in the baseline condition. If network representations simply found any mesh to bi similar to the target, then we would expect the distribution of similarities for random meshes to overlap with that for the reconstructions (seeds). Instead, we find that the models show variable behavior where VGG similarity values are more or less between [0.777,0.020][-0.777,0.020] and ResNet-SIN with very narrow similarities between $[0.997,0.998]$.
Implications
Taken together, these results highlight that the ability of MRD to find metamers is representation-dependent: some feature spaces admit large equivalence classes (making metamers relatively easy to find under our optimization), while others impose stricter constraints that our current reconstruction procedure cannot satisfy. This has two consequences: (1) The baseline run similarity levels (used to define the metamer criterion) must be performed per-representation and per scene. (2) RSA provides an important diagnostic: reconstructions that preserve relational geometry but not pointwise alignment may still be useful in tasks that depend only on representational geometry, whereas pointwise-aligned metamers are necessary when exact code recovery is required.
5. Discussion
We evaluated MRD’s ability to reconstruct physical scene parameters given the information contained in various latent spaces of neural networks. A reconstructed scene is considered a model metamer if the optimization process yields a similarity score equal to or higher than the baseline similarity of the reference scene (optimized against pixel-based loss).
Material Reconstruction.
Across all material experiments, MRD achieved metamerism in a substantial subset of scenes, with many networks reaching or closely approaching their baseline similarities (Table 3). The LPIPS perceptual metric was particularly effective for materials, consistently producing metamers across nearly all tested BSDF classes (Aurora, Brushed Metal, Diffuse, Translucent) with top similarity scores frequently exceeding $0.95$ and minimal deviations from the baseline run. Similarly, VGG and CLIP demonstrated strong performance for several materials, though CLIP exhibited higher variability: certain material–lighting conditions achieved exceptionally high RSA values (e.g., Aurora CLIP, RSA =1.0=1.0) despite moderately lower absolute similarity scores. Although a portion of the BSDF experiments did not strictly meet the metamer criterion, many of these failures were marginal. Networks such as ResNet-SIN often produced reconstructions with extremely small similarity gaps (e.g., $\Delta_{\text{sim}}<0.02$ for brushed metal and translucent materials), indicating that the optimized BSDFs were nearly representationally identical to the baseline. These marginal failures suggest that small amounts of optimization noise or local irregularities in the BSDF parameter landscape, rather than representational mismatch, prevented some reconstructions from formally qualifying as metamers. Overall, material reconstruction appears robust across networks, with LPIPS and VGG providing the highest similarities.
Shape Reconstruction.
Shape reconstruction was more challenging overall, with far fewer scenes meeting our metamer threshold (Table 4). While LPIPS remained a strong metric, achieving metamers across several of the evaluated object classes (Dog, Dragon, Lion Statue, Suzanne), shape recovery generally required more iterations and exhibited greater network-dependent variability. VGG also succeeded in several cases (e.g., Suzanne Garden VGG), reflecting its sensitivity to mid-level image structure that is strongly influenced by shape. However, CLIP and ResNet variants showed more variable results for geometry: in many conditions, similarity fell just short of baseline despite the reconstructions visually matching the target geometry. These marginal failures are particularly prominent for ResNet and ResNet-SIN, where several experiments achieved exceptionally high absolute similarity (often >0.97>0.97) but missed the metameric threshold by only a small margin (e.g., $\Delta_{\text{sim}}\approx 0.016$$0.025$ for Dragon scenes). Such small differences indicate that although the reconstructed shapes did not strictly surpass the baseline similarity, their latent representations were almost identical to the baseline. These results suggest that shape reconstruction is more susceptible to optimization noise and local minima in the high-dimensional geometry space, but that the method nonetheless produces high-quality latents that are very close to metameric for several networks, even though the visual reconstructions often yield “anamorphous blobs" (as seen in Figure 5).
Why Material Reconstruction Outperforms Shape Reconstruction
We observe consistently higher reconstruction similarity and more stable convergence for material (BSDF) reconstruction compared to shape reconstruction. This difference can be explained by several factors. (1) Material variation occupies a substantially lower-dimensional and more structured subspace than shape. Changes in BRDF parameters such as roughness, metallicity, or anisotropy produce smooth, predictable modifications in shading and reflections, which map more linearly into the model’s latent space. In contrast, shape variation is inherently high-dimensional: small geometric perturbations can induce large, non-linear changes in silhouette, occlusion, and local shading, making the latent response more complex and less predictable. (2) Material edits affect the image globally, providing spatially dense and coherent gradients during optimization, whereas shape edits often result in sparse or discontinuous gradients that slow convergence and increase the likelihood of suboptimal solutions. (4) Modern vision networks have been shown to encode texture and shading statistics more explicitly than geometric information (Geirhos et al., 2019); as a result, latent embeddings may be naturally more sensitive to material differences than to shape differences. (5) The optimization landscape for materials is smoother, with fewer local minima, and the baseline material embeddings exhibit lower intra-class variance. Together, these factors make it significantly easier for the optimizer to match the baseline similarity for materials, leading to improved reconstruction quality and more reliable identification of material metamers.
RSA
In addition to hypersphere similarity, we also computed RSA scores to assess the structural correspondence between the model’s internal representations and the generated output (for models in which the latent was used instead of a score). Significant correlations ($p<0.05$) were observed in 28 experiments. Interestingly, high representational alignment did not always guarantee metamerism. For instance, Dog Garden ResNet achieved a strong RSA correlation ($0.621$, Sig: Yes) but failed to reach metamerism ($\Delta_{\text{sim}}=0.080$), implying that the geometry of the manifold was preserved even if the exact point on the manifold was not reached.
Physically Based Light Transport vs. Rasterization
Recently, Elumalai et al. (Elumalai et al., 2025) optimized a mesh to obtain maximally exciting inputs for a Robust ResNet’s activation using a differentiable rasterizer reconstucting. In their work, they successfully use differentiable rendering to optimize against a network activation objective, yet, rasterization lacks physicality as it is bound to approximations of global illumination and/or screen-space techniques. In contrast, MRD reconstructs latent representations using a differentiable path tracer, thereby accounting for physically based light transport and simulating light propagation in a 3D scene.
Concepts in higher layers
Finally, we note that most of the qualitative visualizations we presented here imply that a successful reconstruction is one that appears subjectively similar to the target image. However, depending on the model and the representational layer(s) used in the reconstruction, this is not necessarily the only possible “success”. Consider that Feather and colleagues (Feather et al., 2023) demonstrated that matching the activations of early layers of a convolutional neural network can produce similar images to the target, whereas later layers exhibit images that appear essentially meaningless to humans. We speculate that MRD, due to its physical basis and constraints, may allow researchers to gain insight into the nature of the semantic-level representation of a given network: what features does the model learn to become invariant to in representing a higher-level concept? Consider that asking a human to visualize a “dragon" would likely elicit rich shape-based descriptions. A person would not think of exactly the target dragon, but of other forms that other humans would also consider to be a dragon. We assert that, were it possible to run MRD on a human, the result would be a series of dragon-like meshes that other observers would likewise classify as “dragons". Not all of the same dragon, and not all in the same pose, but dragons nonetheless. If a network existed that had a human-like visual semantic representation of dragons (including shape-based descriptors), then we might expect MRD to reconstruct a family of dragon-like shapes. One interpretation of our present results is that such a network currently does not exist.
5.1. Future work
While our experiments are extensive, we intentionally chose simple scenes to demonstrate the utility of MRD as a proof-of-concept. Future work could study other environmental aspects (camera position, lighting) and also combinations of scene parameters together. Furthermore, more computationally expensive models can be probed. PBRD is an active field in computer graphics, and advances in PBDR (such as Many Worlds Rendering (Zhang et al., 2024)) will help to improve future iterations of MRD.
Our work has additional technical limitations. First, probing larger NNs is non-trivial; we require sharding the model across multiple GPUs. When Mitsuba is involved, inter-process communication is required to send gradients between devices so we can run the backward pass across GPUs. Second, we did not mask the environment maps during the optimization process; therefore, a strong cue for the reconstruction might be anchor points around the silhouette of a shape. Third, rendered images may represent an out-of-distribution problem for the models we tested; fine-tuning on renderings may improve results. Finally, because we test our method only on rendered images, one reason the models perform so poorly in shape reconstruction may be that they are not, in general, trained on our scenes. Our scenes view objects from unusual views below the object, while in ImageNet, objects are primarily viewed from above or the side.
6. Conclusion
We present MRD (metamers rendered differentiably), a novel approach to reasoning about image-computable models in physically grounded 3D space. Our method can reconstruct model metamers, enabling us to determine whether a model is invariant or responsive to a given physical scene parameter (such as material, or shape). MRD demonstrates a robust capacity for generating metameric scenes, particularly when guided by perceptual losses. For standard architectural backbones, the results indicate a soft success, where the reconstructions consistently fall within a tight margin of the baseline, validating the efficacy of the proposed latent space reconstruction technique. As the first work in this direction, there are significant limitations and opportunities for improvement.
Author contributions
The original idea to apply PBDR as an evaluation method to understand learned visual representations was developed jointly by BB and TSAW. BB performed all implementation and analysis. BB and TSAW jointly wrote and edited the paper.
Acknowledgments
We thank Joshua Martin for feedback on the manuscript. We further thank the following artists and authors for providing the assets for non-commercial purposes: Delatronic (Material Dragon), Stanford’s Dragon model (Shape Dragon), Rigsters’ Lion crushing a serpent (Lion Statue), and Blender’s Suzanne. For the Hallstatt envmap, we thank Bernhard Vogl. The other three environment maps were available on HDRI Haven. The Aurora and Brushed Metal materials were measured and provided by Dupey and Jakob (Dupuy and Jakob, 2018).
Disclosure of Machine Learning Tools Use
In the preparation of this manuscript, we use large language models (LLM) for orthographic purposes, such as the correction of spelling and grammar and minor rephrasing.
Funding
Funded by the European Union (ERC, SEGMENT, 101086774). Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. This work was additionally supported by the Deutsche Forschungsgemeinschaft (German Research Foundation, DFG) under Germany’s Excellence Strategy (EXC 3066/1 “The Adaptive Mind”, Project No. 533717223).
Appendix A Supplementary Material
We provide more qualitative results for both material and shape reconstructions as seen in Figures [A7 - A11].
Figure A7. Reconstruction of a translucent, textured material using DINO and LPIPS. DINO manages to find suitable parameters to make the model translucent, however, it fails to reconstruct the texture. LPIPS is closer to the baseline but the texture reconstruction appears more noisy.
Figure A8. Aurora material reconstruction with CLIP and VGG. Both models show a suitable reconstruction of a roughconductor, but differ in the amount of noise in the texture.
Figure A9. Reconstruction of the dog mesh in the garden environment map using VGG and DINO.
Figure A10. Reconstruction of the lion statue mesh in the Hallstatt environment map using LPIPS and CLIP.
Figure A11. Reconstruction of the Suzanne mesh in the skybox environment map using VGG and ResNet-SIN.